Grade 12Physical Chemistry

The first ionisation energy of potassium is 100kcal mol-1.Calculate the lowest possible frequency of light that can ionise a potassium atom.

Nitin
8 Years agoGrade 12

1 Answer

Mahima Kanawat
8 Years ago
Dear student .....To ionise a K atom we require minimum 100 kcal ( 100×10³×4.2 J) energy.......As we know E=hv.....42×10⁴=6.6×10³⁴×v.....v=6.3×10^-30Hz......With Regards .....Mahima..(Askiitians forum expert )